Can you please help me with this?
determine which ordered pairs satisfy the given eguation.
4x - y = 4 (0,-4) (1,0) (2,-3)
I need to know how it is done, I never had this in school. Can you please show me how to do it? Thank You!

Answer by checkley77(12844)

4x - y = 4 (0,-4) (1,0) (2,-3)
The way to prove if these points are on this line.
Replace the x & y values & solve the equation.
4*0-(-4)=4
0+4=4
4=4 (0,-4) is on this line.
4*1-0=4
4-0=4
4=4 (1,0) is also on this line.
2*2-(-3)=4
4+3=4
7=4 not a solution thus this point is NOT on the line.